OpenStudy (anonymous):

Simplify square root of 3 times square root of 21.

OpenStudy (maddy1251):

@Skittles8082 need help?

OpenStudy (maddy1251):

Also, is this\[\sqrt{3} \times \sqrt{21}\]

OpenStudy (mokana219):

Simple, when multiplying square roots are squishy, all you have to do is multiply 3 by 21 and then simplify that answer by factoring.